No.
What is the probability of flipping a coin 4 times and getting heads each time?
The probability on the first flip is 50% .
The probability on the 2nd flip is 50% .
The probability on the 3rd flip is 50% .
The probability on the 4th flip is 50% .
The probability of 4 heads is (50% x 50% x 50% x 50%) = (0.5)4 = 1/16 = 6.25%
Assess hazards is the step in the composite risk management process that is focused on determining the probability and severity of a hazard occurring.
5/12
If it's wrong don't mind me, i'm not even in middle school yet i'm just smart
About 2/3 of the time.
Anyone who is going to answer this will need to know what choices there are for the answers. Without knowing this they will not be able to know which cause something dangerous to happen.
Comparision between binomial plus normal plus hypergeometric distribution?
The binomial distribution is a discrete probability distribution which describes the number of successes in a sequence of draws from a finite population, with replacement. The hypergeometric distribution is similar except that it deals with draws without replacement. For sufficiently large populations the Normal distribution is a good approximation for both.
a is selected by 63 out of 287 - a rate of 63/287.
b is selected at a rate of 23/301.
Selection rate for a/selection rate for b = (63/287)/(23/301) = 0.2195/0.0764 = 2.8727
Ie a is likely to be selected 2.8727 times as often as b, or a is 1.8727 times more likely to be selected.
If you toss a penny how do you think the coin will land heads or tails explain?
there is even chance of having heads or tails since there is only 1 head and 1 tail on a penny.
What is the probability of flipping NO heads when you flip 3 identical coins?
The chance of not flipping a head in each instance is 1/2. You need that to happen three times.
1/2 x 1/2 x 1/2 = 1/8
So there is a 1 in 8 chance of getting no heads from 3 coin flips.
What is the probability of rolling a sum which is a multiple of five?
There are 36 permutations of rolling two standard dice. Of them, four (1+4, 2+3, 3+2, and 4+1) sum to five, and three (6+4, 5+5, and 4+6) sum to ten. The probability, then of rolling a multiple of five is 7 in 36, or about 0.1944.
How many ways are there to arrange the letters in the word ADDRESS if the two Ds must be together?
The permutations of the letters ADDRESS if the two D's must be together is the same as the permutations of the letters ADRESS, which is 6 factorial, or 720, divided by 2, to compensate for the two S's, which means that the number of distinct permutations of the letters ADDRESS, where the two DD's must be together is 360.
How many card that is not a diamond face card?
There are 52 cards in a deck. 52cards/4 suits = 13. 13x3=39. There are 39 cards in a deck that are not diamond.
When apply geomatric distribution?
When you want to find the probability of something happening for the first time while repeating an experiment, you use the geometric distribution. For example, you are throwing a six-faced die and you are expecting the number '6' to be thrown for the first time.
Lucy and her friends each have seven cards, as one third of 21 is 7.
A definition and an example of multiple representation?
it's basically higher order thinking ex) 3+4-2*4dived by 8= e